OpenStudy (anonymous):

-18=p/12 please help

OpenStudy (lucaz):

multiply both sides of the equation by 12 and you will find the value of p

OpenStudy (lucaz):

\[\large{(12)\times -18=\frac{p}{\cancel{12}}\times (\cancel{12})}\]